TV advertising agencies face increasing challenges in reaching audience members because viewing TV programs via digital streaming is gaining in popularity. A poll reported that 58% of 2341 American adults surveyed said they have watched digitally streamed TV programming on some type of device.

(a)

Calculate and interpret a confidence interval at the 99% confidence level for the proportion of all adult Americans who watched streamed programming up to that point in time. (Round your answers to three decimal places.)

  ,

Interpret the resulting interval.

We are 99% confident that this interval contains the true population proportion.We are 99% confident that this interval does not contain the true population proportion.    We are 99% confident that the true population proportion lies above this interval.We are 99% confident that the true population proportion lies below this interval.

(b)

What sample size would be required for the width of a 99% CI to be at most 0.04 irrespective of the value of ? (Round your answer up to the nearest integer.)
Answer #1

The statistical software output for this problem is:

1596230073173_image.png

From above output:

a) 99% confidence interval:

(0.554, 0.606)

Interpretation: We are 99% confident that this interval contains the true population proportion.

b) Given,

p = 0.58, q = 1 - 0.58 = 0.42

E = 0.04/2 = 0.02

For 99% confidence, z = 2.576

Hence,

Sample size required

n = (0.58)(0.42)(\frac{2.576}{0.02})^2

n = 4042